a triangle has a hypotenuse of length 15 and a leg length of 8. What is the area of the triangle
Answer by mananth(16946)

In a right triangle
hypotenuse^2 = leg1^2 +leg2^2
hypotenuse^2 - leg1^2 =leg2^2
15^2-8^2= leg2^2
225-64= leg2^2
161= leg2^2
leg2 = sqrt161
..
Area of right triangle = 1/2 * leg1 * leg2
A= 1/2 * 8*sqrt161
Area = 50.75 sq units